# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

### v2

@@ 4. Conjectures on algebraic cycles
We now review the conjectures on algebraic cycles we will be concerned with, and give some sufficient conditions for them.
@@ 4.1. The generalized conjectures of Tate and Beilinson
For $k$ a field, $X$ a scheme with finite-type, separated structure morphism $f:X\to\operatorname{Spec}k$, and $i,j\in{\mathbb{Z}}$, we define the $\ell$-adic étale homology of $X$ as in [Jannsen1990, Example 6.7], via
@@
Define the (rational) motivic homology and cohomology of $X$ as in [Jannsen1990, 5.12 and Example 6.12], via
@@
Here $K_{m}(X)$ denotes the $m$-th Quillen $K$-theory of vector bundles on $X$, and $K_{m}(X)^{(r)}$ denotes the simultaneous eigenspace of $K_{m}(X)\otimes_{\mathbb{Z}}{\mathbb{Q}}$ for the Adams operations $\{\psi^{n}:n\in{\mathbb{Z}}-\{0\}\}$, on which $\psi^{n}$ acts as multiplication by $n^{r}$. The definition of $K^{\prime}_{m}(X)^{(r)}$ is the same, except using the $K$-theory of coherent sheaves on $X$ instead of vector bundles. See [Soule1985, 1.5 and 7.2] for the definition of the Adams operations.
For context, we note that for a smooth variety $X$ over $k$, and $i$, $j\in{\mathbb{Z}}$, there is a a canonical isomorphism
@@
cf. [Bloch1986, page 269(vii)] and [Jannsen1990, Example 6.12]. On the right is Bloch’s higher Chow group (with $\mathbb{Q}$}-coefficients). For such an $X$, we thus let $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ denote $\operatorname{CH}^{j}(X,2j-i)_{\mathbb{Z}}$. The more modern definitions of integral motivic cohomology and Borel–Moore homology are not in terms of higher Chow groups or $K$-theory [MVW2006, Definitions 3.4 and 16.20], but this comparatively classical approach will suffice for us.
In any case, we will not be too preoccupied with the precise definition of motivic (co)homology. Both étale and motivic (co)homology form twisted Poincaré duality theories in the sense of [Jannsen1990, Definition 6.1] [^1] There is a caveat in the case of motivic (co)homology, which forms an “absolute” twisted Poincaré duality theory, rather than a relative one. See [Jannsen1990, Example 6.12]., so that we have the following purely formal tools for working with them. Out of an abundance of caution, in the remainder of this section we work in the category of quasi-projective schemes over the base field, unless otherwise noted [^2] We could possibly get away with weaker hypotheses, but [Soule1985] as stated works in this generality..
@@
    For $0\leq i\leq\dim X$, there is a canonical isomorphism $H_{2i}^{\mathcal{M}}(X,{\mathbb{Q}}(i))\cong\operatorname{CH}_{i}(X)$ [Jannsen1990, Lemma 6.12.4]. For $i,j\in{\mathbb{Z}}$, there is a regulator map
@@
In terms of motivic and étale homology, Jannsen generalized conjectures of Tate [Tate1994, Conjecture $T^{j}(X)$] and Beilinson [Beilinson1987, page 3] to arbitrary finite-type separated schemes over ${\mathbb{F}}_{q}$.
Conjecture 4.2 ($\operatorname{TB}_{i,j}(X)$: [Jannsen2007, Conjecture 4.14]).
@@
Here we define a finite-dimensional $G_{{\mathbb{F}}_{q}}$-representation $V$ to be 1-semisimple if 1 is a semisimple eigenvalue of the geometric Frobenius $\operatorname{Frob}_{q}\in G_{{\mathbb{F}}_{q}}$ acting on $V$. We denote the truth value of the conjecture for a given $(X,i,j)$ by $\operatorname{TB}_{i,j}(X)$. By $\operatorname{TB}(X)$ we mean the truth value of the conjecture for a given $X$ and all $i,j\in{\mathbb{Z}}$. When $X$ is of pure dimension $d$, we define $\operatorname{TB}^{i,j}(X)=\operatorname{TB}_{2d-i,d-j}(X)$ (so the notation is compatible with Poincaré duality for a smooth $X$). The conjecture $\operatorname{TB}$, taken for all smooth projective varieties over ${\mathbb{F}}_{q}$, simultaneously generalizes the Tate–Beilinson and Parshin conjectures [Kahn2005, Conjecture 51 and Definition 55]. See loc. cit. for an overview of the general consequences of these conjectures.
@@
Proposition 4.3 ([Broe2025, Proof of Lemma 4.4]).
@@
    If $Z$ is a closed subscheme of $X$ with open complement $U$, and $\operatorname{TB}$ holds for any two of $X$, $Z$ and $U$, then it also holds for the third one [Jannsen1990, Theorem 12.7 b)].
@@
    If $X$ is a smooth projective variety with finite-dimensional motive, the Tate conjecture holds for $X$, and for all $i$, $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple, then $\operatorname{TB}(X)$ holds ([Jannsen2007, Theorem 4.12]).
@@
    If $X$ is a smooth projective variety, rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$, and the Tate conjecture holds for cycles of codimension $j$ in $X$, then $\operatorname{TB}^{2j,j}(X)$ holds [Tate1994, Theorem 2.9].
@@
The left arrow is injective by the projection formula (proposition 4.1(iv)), and the bottom arrow is injective by $\operatorname{TB}_{i,j}(X_{L})$, so the the top arrow is injective.
@@
We remark that when $X$ is a smooth projective variety over ${\mathbb{F}}_{q}$, rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$ when $j=1$ (by the theory of the Picard variety and [Broe2025, Lemma 2.3]) or when $j=\dim X$ (by the theory of the Albanese variety and proposition 2.4).
@@
Using proposition 4.5(ii), we may assume that $X$ has a rational point. The claim then follows from [EKM2008, Proposition 70.1] and induction on the dimension of $X$. ∎
@@
Claim (i) is clear from proposition 4.5 and the semisimplicity of the $\ell$-adic cohomology of abelian varieties over finite fields, which follows from [Tate1966, page 138] and Chevalley’s theorem [Maculan2018, Theorem 1].
In the situation of (ii), by [Vial2017, Theorem 3.11 and Proposition 3.13], there exists a direct sum decomposition
@@
where $Q$ is a direct summand of $\mathfrak{h}(Y)\oplus\mathfrak{h}(C_{1})\oplus...\oplus\mathfrak{h}(C_{m})$ for some curves $C_{1},...,C_{m}$, and $R$ is a direct summand of $\mathfrak{h}(S_{1})\oplus...\oplus\mathfrak{h}(S_{n})$ for some varieties $S_{1},...,S_{n}$ of dimension $\leq 2$ [^3] Note that [Vial2017] uses a “covariant” convention for motives, and the opposite of our sign convention for Tate twists. Also, its definition of “variety” does not require irreducibility.. Theorem 3.12 of loc. cit. additionally implies that the motive of $X$ is finite-dimensional. Using proposition 4.5(ii), we are free to extend scalars and assume that all $C_{i}$ and $S_{j}$ are geometrically connected. It then follows from the decomposition of the motive that $X$ satisfies the Tate conjecture for divisors, and that $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple for $i\leq 1$. The full Tate conjecture for $X$, and the 1-semisimplicity of its cohomology in higher degrees, then follows from the hard Lefschetz theorem [Deligne1980, Théorème (4.1.1)]. We conclude by proposition 4.5(iii). ∎
@@
The conjecture $\operatorname{TB}(X)$ holds when $X\subset\mathbb{P}^{n+1}$ is a Fermat hypersurface of degree $m$, and $p=\operatorname{char}({\mathbb{F}}_{q})$ satisfies $p^{r}\equiv-1\pmod{m}$ for some $r$.
@@
Such an $X$ has motive of abelian type, and satisfies the Tate conjecture [Tate1965, page 102]. The claim thus follows from corollary 4.7. ∎
@@
The claim is clear for $\dim X=1$, by proposition 4.5. Assume $\dim X>1$. By the same proposition, we may pass to a finite extension and assume that $X$ has a rational point. Then by [Kollar2002, Theorem 1], $X$ is unirational, so corollary 4.7 applies. ∎
@@
By corollary 4.9 we may assume $d\geq 4$. The claim then follows from combining proposition 4.5(iv) with the results of section 2 on Chow groups of cubic hypersurfaces. Specifically, when $j=1$, we use the Tate conjecture for divisors on $X$; when $j=2$, we use the agreement of rational and numerical equivalence on $\operatorname{CH}^{2}(X)$ (proposition 2.2) and the Tate conjecture in codimension 2 for $X$ (proposition 2.7); when j=d1j=d-1 we argue similarly, via corollary 2.5 and proposition 3.3, and when $j=d$ we use the finiteness of the Albanese kernel (proposition 2.4). ∎
We note that Rapoport proved in general that the $\ell$-adic cohomology of a smooth cubic fourfold over ${\mathbb{F}}_{q}$ is a semisimple Galois representation ([Rapoport1972]; see also [Zarhin2004, page 1]).
Using Voevosky’s proof of the Bloch–Kato conjecture [Voe2011], one can get some limited control over the integral motivic cohomology of varieties satisfying the conjecture $\operatorname{TB}$.
@@
This follows from the argument of Corollaries 69 and 70 of [Kahn2005]. We will fill in some details for the reader who wishes to check them, referring to the statements of loc. cit. as numbered therein. The crucial point is to observe, from the equivalent formulation of the Bloch–Kato conjecture in Lemma 18, that $H^{i}_{\mathcal{M}}(X,\mathbb{Z}(j))$ injects into $H^{i}_{\operatorname{\acute{e}t}}(X,\mathbb{Z}(j))$ for $i\leq j+2$. Here the latter group is étale motivic cohomology. Now use the finite generation of the Weil-étale cohomology of $X$ (proved in Theorem 67) and the long exact sequence of (4.39). Recalling that $H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))\cong H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Q}}(j))$ (Theorem 14), one finds that $H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Z}}(j))$ can only fail to be finitely generated when $i=2j+2$. From this one concludes that $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ is finitely generated whenever $i\leq j+2$ and $j>0$. The lower bound on $j$ is removed through the straightforward observation that when $j\leq 0$, one has
@@ 4.2. The Beilinson–Bloch conjecture
This conjecture was independently formulated over number fields by Beilinson [Beilinson1987, Conjecture 5.0] and Bloch [Bloch1984, page 94] (and both attributed it to Swinnerton-Dyer, who seems not to have published it). We write $\operatorname{BB}^{i}(X)$ for the truth value of the conjecture for a given $X$ and $i$. Sometimes included along with conjecture 4.12 are the statements that the Chow group with integer coefficients $\operatorname{CH}^{i}(X)_{\mathbb{Z}}$ is finitely generated (the “motivic Bass conjecture”) [^4] Note that Kahn expresses some doubt about the analogous assertion over finite fields in [Kahn2005, Section 4.9.1]., and that the Abel–Jacobi map
@@
is injective [Jannsen1990, Conjecture 9.15].
Little is known in general about the Beilinson–Bloch conjecture, and we will not attempt more than a very limited recollection of existing results (see [LTXZZ2025, Remark 1.3] for some additional background). When $k={\mathbb{Q}}$ or when $\operatorname{char}(k)>0$, the conjecture is known for elliptic curves over $k$ of analytic rank at most one: this claim over $\mathbb{Q}$} follows from [GZ1986], [Koly1988], [Wiles1995], and [BCDT01], and over function fields from [Qiu2022, Theorems 12.1.2 and 12.3.1]. When $k$ is a number field and $i>1$, partial evidence has come from the study of Ceresa cycles (as in [Bloch1984]). The conjecture $\operatorname{TB}$, taken for all varieties over finite fields, is known to imply the Beilinson–Bloch conjecture in positive characteristic ([Jannsen1990, Chapter 12], [Broe2025, Corollary 5.6]), and one can use this approach to prove the conjecture unconditionally in special cases (e.g. in the cases mentioned in the introduction, and in [Ulmer2011, Theorem 3.9.1]). In particular, we have the following criterion, which will be our main tool in proving cases of the Beilinson–Bloch conjecture later in the paper.
Proposition 4.13 ([Broe2025, Remark 5.3 and Corollary 5.6]).
@@
surjects onto the Galois-invariant subspace, its target $H_{2d-2j}((X_{Z})_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(d-j))$ is 1-semisimple, and $\operatorname{TB}^{2j,j}(X)$ holds, then $BB^{j}(X_{\eta})$ holds, and the Abel–Jacobi map
$\operatorname{CH}^{j}(X_{\eta})_{0}\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\to H^{1}_{\operatorname{cont}}(G_{k(C)},H^{2j-1}(X_{\overline{k(C)}},{\mathbb{Q}}_{\ell})))$
is bijective. Here $H^{1}_{\operatorname{cont}}$ denotes continuous group cohomology.
We remark that it would follow from the Beilinson–Bloch conjecture that the Albanese kernel of any smooth, projective, geometrically connected variety over a global field is torsion [Beilinson1987, Lemma 5.1].
Tne Beilinson–Bloch conjecture, in the form stated here, is closely related to conjectures of Beilinson, Bloch and Murre describing filtrations on the Chow groups of smooth projective varieties over arbitrary fields. We refer to Jannsen’s article [Jannsen1994] for a wonderful overview of some of the ideas behind and connections between the conjectures.

### v3

@@ 4. Conjectures on algebraic cycles
We now review the conjectures on algebraic cycles we will be concerned with, and give some sufficient conditions for them. Proposition 4.14 shows that a sufficiently strong form of the Tate conjecture over finite fields (conjecture 4.2) implies the Beilinson–Bloch conjecture over function fields, so our main results will hinge on verifying cases of the former conjecture.
@@ 4.1. The generalized conjectures of Tate and Beilinson
For $k$ a field, $X$ a scheme with finite-type, separated structure morphism $f:X\to\operatorname{Spec}k$, and $i,j\in{\mathbb{Z}}$, we define the $\ell$-adic étale homology of $X$ as in [47, Example 6.7], via
@@
Define the (rational) motivic homology and cohomology of $X$ as in [47, 5.12 and Example 6.12], via
@@
Here $K_{m}(X)$ denotes the $m$-th Quillen $K$-theory of vector bundles on $X$, and $K_{m}(X)^{(r)}$ denotes the simultaneous eigenspace of $K_{m}(X)\otimes_{\mathbb{Z}}{\mathbb{Q}}$ for the Adams operations $\{\psi^{n}:n\in{\mathbb{Z}}-\{0\}\}$, on which $\psi^{n}$ acts as multiplication by $n^{r}$. The definition of $K^{\prime}_{m}(X)^{(r)}$ is the same, except using the $K$-theory of coherent sheaves on $X$ instead of vector bundles. See [90, 1.5 and 7.2] for the definition of the Adams operations.
For context, we note that for a smooth variety $X$ over $k$, and $i$, $j\in{\mathbb{Z}}$, there is a canonical isomorphism
@@
cf. [10, page 269(vii)] and [47, Example 6.12]. On the right is Bloch’s higher Chow group (with $\mathbb{Q}$}-coefficients). For such an $X$, we thus let $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ denote $\operatorname{CH}^{j}(X,2j-i)_{\mathbb{Z}}$. The more modern definitions of integral motivic cohomology and Borel–Moore homology are not in terms of higher Chow groups or $K$-theory [74, Definitions 3.4 and 16.20], but this comparatively classical approach will suffice for us. Corollary 5.3 of [67] shows that $\mathbb{Q}$}-linear correspondences of smooth projective varieties act on higher Chow groups with $\mathbb{Q}$}-coefficients, so $X\mapsto H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))$ extends to a covariant functor on $\mathcal{M}(k)$.
In any case, we will not be too preoccupied with the precise definition of motivic (co)homology. Both étale and motivic (co)homology form twisted Poincaré duality theories in the sense of [47, Definition 6.1] [^2] There is a caveat in the case of motivic (co)homology, which forms an “absolute” twisted Poincaré duality theory, rather than a relative one. See [47, Example 6.12]., so that we have the following purely formal tools for working with them. Out of an abundance of caution, in the remainder of this section we work in the category of quasi-projective schemes over the base field, unless otherwise noted [^3] We could possibly get away with weaker hypotheses, but [90] as stated works in this generality..
@@
    For $0\leq i\leq\dim X$, there is a canonical isomorphism $H_{2i}^{\mathcal{M}}(X,{\mathbb{Q}}(i))\cong\operatorname{CH}_{i}(X)$ [47, Lemma 6.12.4]. For $i,j\in{\mathbb{Z}}$, there is a regulator map
@@
In terms of motivic and étale homology, Jannsen generalized conjectures of Tate [94, Conjecture $T^{j}(X)$] and Beilinson [8, page 3] to arbitrary finite-type separated schemes over ${\mathbb{F}}_{q}$.
Conjecture 4.2 ($\operatorname{TB}_{i,j}(X)$: [47, Conjectures 12.4a and 12.6b]).
@@
Here we define a finite-dimensional $G_{{\mathbb{F}}_{q}}$-representation $V$ to be 1-semisimple if 1 is a semisimple eigenvalue of the geometric Frobenius $\operatorname{Frob}_{q}\in G_{{\mathbb{F}}_{q}}$ acting on $V$. We denote the truth value of the conjecture for a given $(X,i,j)$ by $\operatorname{TB}_{i,j}(X)$. By $\operatorname{TB}(X)$ we mean the truth value of the conjecture for a given $X$ and all $i,j\in{\mathbb{Z}}$. When $X$ is of pure dimension $d$, we define $\operatorname{TB}^{i,j}(X)=\operatorname{TB}_{2d-i,d-j}(X)$ (so the notation is compatible with Poincaré duality for a smooth $X$). The conjecture $\operatorname{TB}$, taken for all smooth projective varieties over ${\mathbb{F}}_{q}$, simultaneously generalizes the Tate–Beilinson and Parshin conjectures [54, Conjecture 51 and Definition 55]. See loc. cit. for an overview of the general consequences of these conjectures.
@@
Proposition 4.3 ([14, Proposition 3.1 and Lemma 3.6]).
@@
    If $Z$ is a closed subscheme of $X$ with open complement $U$, and $\operatorname{TB}$ holds for any two of $X$, $Z$ and $U$, then it also holds for the third one [47, Theorem 12.7 b)].
@@
    If $X$ is a smooth projective variety with finite-dimensional motive, the Tate conjecture holds for $X$, and for all $j$, $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple, then $\operatorname{TB}(X)$ holds ([50, Corollary 13.5.10]).
@@
    If $X$ is a smooth projective variety, rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$, and the Tate conjecture holds for cycles of codimension $j$ in $X$, then $\operatorname{TB}^{2j,j}(X)$ holds [94, Theorem 2.9].
@@
The left arrow is injective by the projection formula (proposition 4.1(iv)), and the bottom arrow is injective by $\operatorname{TB}_{i,j}(X_{L})$, so the top arrow is injective.
@@
We remark that if $X$ is a smooth projective variety over ${\mathbb{F}}_{q}$, then rational and numerical equivalence coincide on $\operatorname{CH}^{j}(X)$ when $j=1$ (by the theory of the Picard variety and [14, Lemma 2.3]) or when $j=\dim X$ (by the theory of the Albanese variety and proposition 2.5).
@@
Using proposition 4.5(ii), we may assume that $X$ has a rational point. The claim then follows from [32, Proposition 70.1] and induction on the dimension of $X$. ∎
@@
Claim (i) is clear from proposition 4.5(iii) and the semisimplicity of the $\ell$-adic cohomology of abelian varieties over finite fields, which follows from [92, page 138] and Chevalley’s theorem [71, Theorem 1].
In the situation of (ii), by [100, Theorem 3.11 and Proposition 3.13], there exists a direct sum decomposition
@@
where $Q$ is a direct summand of $\mathfrak{h}(Y)\oplus\mathfrak{h}(C_{1})\oplus...\oplus\mathfrak{h}(C_{m})$ for some curves $C_{1},...,C_{m}$, and $R$ is a direct summand of $\mathfrak{h}(S_{1})\oplus...\oplus\mathfrak{h}(S_{n})$ for some varieties $S_{1},...,S_{n}$ of dimension $\leq 2$ [^4] Note that [100] uses a definition of “variety” that does not require irreducibility.. Theorem 3.12 of loc. cit. additionally implies that the motive of $X$ is finite-dimensional. Using proposition 4.5(ii), we are free to extend scalars and assume that all $C_{i}$ and $S_{j}$ are geometrically connected. It then follows from the decomposition of the motive that $X$ satisfies the Tate conjecture for divisors, and that $H^{2j}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(j))$ is 1-semisimple for $j\leq 1$. The full Tate conjecture for $X$, and the 1-semisimplicity of its cohomology in higher degrees, then follows from the hard Lefschetz theorem [29, Théorème (4.1.1)]. We conclude by proposition 4.5(iii). ∎
@@
The conjecture $\operatorname{TB}(X)$ holds when $X\subset\mathbb{P}^{n+1}_{{\mathbb{F}}_{q}}$ is a Fermat hypersurface of degree $m$, and either
  1. (i)
    $p=\operatorname{char}({\mathbb{F}}_{q})$ satisfies $p^{r}\equiv-1\pmod{m}$ for some $r$, or
  2. (ii)
    $n$ is odd.
@@
Such an $X$ has motive of abelian type by [58, Theorem 1]. It also satisfies the Tate conjecture, by [91, page 102] in case (i), and proposition 3.7 in case (ii). The claim thus follows from corollary 4.7. ∎
@@
The claim is clear for $\dim X=1$, by proposition 4.5(iii). Assume that $\dim X>1$. By proposition 4.5(ii), we may pass to a finite extension and assume that $X$ has a rational point. Then by [61, Theorem 1], $X$ is unirational, so corollary 4.7 applies. ∎
@@
By corollary 4.9, we may assume that $d\geq 4$. The claim then follows from combining proposition 4.5(iv) with the results of section 2 on Chow groups of cubic hypersurfaces. Specifically, when $j=1$, we use the Tate conjecture for divisors on $X$; when $j=2$, we use the agreement of rational and numerical equivalence on $\operatorname{CH}^{2}(X)$ (proposition 2.2) and the Tate conjecture in codimension 2 for $X$ (propositions 2.8 and 3.7); when j=d1j=d-1 we argue similarly, via corollary 2.6, and when $j=d$ we use the finiteness of the Albanese kernel (proposition 2.5). ∎
Using the previous proposition, we can show that the Fermat cubic fourfold sits in a larger family of cubic fourfolds satisfying the conjecture $\operatorname{TB}$.
Corollary 4.11.
Assume that $\operatorname{char}({\mathbb{F}}_{q})\neq 3$. Let $X$ be a smooth cubic fourfold over ${\mathbb{F}}_{q}$, defined in $\mathbb{P}^{5}_{{\mathbb{F}}_{q}}$ by an equation of the form
$f(x_{0},...,x_{4})+x_{5}^{3}=0,$
where $f(x_{0},...,x_{4})=0$ defines a smooth cubic threefold in $\mathbb{P}^{4}_{{\mathbb{F}}_{q}}$. Then $\operatorname{TB}(X)$ holds.
Proof.
As noted in remark 3.25, it follows from the main result of [65], and a specialization argument as in the proof of corollary 3.21, that $X$ has finite-dimensional motive. By proposition 4.10, $\operatorname{TB}^{2j,j}(X)$ holds for all $j$, and proposition 4.5(iii) then implies $\operatorname{TB}(X)$. ∎
We note that Rapoport proved in general that the $\ell$-adic cohomology of a smooth cubic fourfold over ${\mathbb{F}}_{q}$ is a semisimple Galois representation ([85]; see also [105, page 45]).
Using Voevodsky’s proof of the Bloch–Kato conjecture [101], one can get some limited control over the integral motivic cohomology of varieties satisfying the conjecture $\operatorname{TB}$.
@@
This follows from the argument of Corollaries 69 and 70 of [54]. We will fill in some details for the reader who wishes to check them, referring to the statements of loc. cit. as numbered therein. The crucial point is to observe, from the equivalent formulation of the Bloch–Kato conjecture in Lemma 18, that $H^{i}_{\mathcal{M}}(X,\mathbb{Z}(j))$ injects into $H^{i}_{\operatorname{\acute{e}t}}(X,\mathbb{Z}(j))$ for $i\leq j+2$. Here the latter group is étale motivic cohomology. Now use the finite generation of the Weil-étale cohomology of $X$ (proved in Theorem 67) and the long exact sequence of (4.39). Recalling that $H^{i}_{\mathcal{M}}(X,{\mathbb{Q}}(j))\cong H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Q}}(j))$ (Theorem 14), one finds that $H^{i}_{\operatorname{\acute{e}t}}(X,{\mathbb{Z}}(j))$ can only fail to be finitely generated when $i=2j+2$. From this one concludes that $H^{i}_{\mathcal{M}}(X,{\mathbb{Z}}(j))$ is finitely generated whenever $i\leq j+2$ and $j>0$. The lower bound on $j$ is removed through the straightforward observation that when $j\leq 0$, one has
@@ 4.2. The Beilinson–Bloch conjecture
This conjecture was independently formulated over number fields by Beilinson [8, Conjecture 5.0] and Bloch [9, page 94] (and both attributed it to Swinnerton-Dyer, who seems not to have published it). We write $\operatorname{BB}^{i}(X)$ for the truth value of the conjecture for a given $X$ and $i$. Sometimes included along with conjecture 4.13 are the statements that the Chow group with integer coefficients $\operatorname{CH}^{i}(X)_{\mathbb{Z}}$ is finitely generated (the “motivic Bass conjecture”) [^5] Note that Kahn expresses some doubt about the analogous assertion over finite fields in [54, Section 4.9.1]., and that the Abel–Jacobi map
@@
is injective [47, Conjecture 9.15]. Here $H^{1}_{\operatorname{cont}}$ denotes continuous group cohomology.
Little is known in general about the Beilinson–Bloch conjecture, and we will not attempt more than a very limited recollection of existing results (see [69, Remark 1.3] for some additional background). When $k={\mathbb{Q}}$ or when $\operatorname{char}(k)>0$, the conjecture is known for elliptic curves over $k$ of analytic rank at most one: this claim over $\mathbb{Q}$} follows from [40], [62], [102], and [13], and over function fields from [84, Theorems 12.1.2 and 12.3.1]. When $k$ is a number field and $i>1$, partial evidence has come from the study of Ceresa cycles (as in [9]). The conjecture $\operatorname{TB}$, taken for all varieties over finite fields, is known to imply the Beilinson–Bloch conjecture in positive characteristic ([47, Chapter 12], [14, Corollary 4.15]), and one can use this approach to prove the conjecture unconditionally in special cases (e.g. in the cases mentioned in the introduction, and in [96, Theorem 3.9.1]). In particular, we have the following criterion, which will be our main tool in proving cases of the Beilinson–Bloch conjecture later in the paper.
Proposition 4.14 ([14, Remark 4.14 and Corollary 4.15]).
@@
surjects onto the Galois-invariant subspace, its target $H_{2d-2j}((X_{Z})_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(d-j))$ is 1-semisimple, and $\operatorname{TB}^{2j,j}(X)$ holds, then $\operatorname{BB}^{j}(X_{\eta})$ holds, and the Abel–Jacobi map
(4.3) $\operatorname{CH}^{j}(X_{\eta})_{0}\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\to H^{1}_{\operatorname{cont}}(G_{k(C)},H^{2j-1}(X_{\overline{k(C)}},{\mathbb{Q}}_{\ell}(j)))$
is bijective.
We remark that it would follow from the Beilinson–Bloch conjecture that the Albanese kernel of any smooth, projective, geometrically connected variety over a global field is torsion [8, Lemma 5.1].
The Beilinson–Bloch conjecture, in the form stated here, is closely related to conjectures of Beilinson, Bloch and Murre describing filtrations on the Chow groups of smooth projective varieties over arbitrary fields. We refer to Jannsen’s article [49] for a wonderful overview of some of the ideas behind and connections between the conjectures.